(include formula)

You currently have $2,500 (Present Value) in an account that has an interest rate of 5% per year compounded monthly (12 times per year). You want to withdraw all your money when it reaches $4,750 (Future Value). In how many years will you be able to withdraw all your money?

The number of years is :
Round your answer to 1 decimal place.
Answer #1

Ans 12.8 years

FV = Future Value
PV = Present Value
r = rate of interest
n= no of period
FV/ PV = (1 + r )^n
4750 / 2500 = (1 + 5%/12)^n
1.9 = 1.00416666666667 ^ n
n = 154 months
n = 12.8 years